OpenStudy (anonymous):

how do you factor x²-xy-2y² ?

OpenStudy (anonymous):

x²-xy-2y² =x^2 -2xy +xy -2y^2 =x(x-2y) +y(x-2y) =(x-2y)(x+y)

OpenStudy (anonymous):

thank you! (: